OpenStudy (anonymous):

x + 3y = 7 x - 3y = 1 Solve?

OpenStudy (texaschic101):

do you know how to add them ?

OpenStudy (anonymous):

im in 9th and we havent done this yet but its on my homework

OpenStudy (texaschic101):

x + 3y = 7 x - 3y = 1 -------------add 2x = 8 -- divide both sides by 2 x = 4 now that we know x, we can sub 8 in for x in either of the original equations x - 3y = 1 4 - 3y = 1 -3y = 1 - 4 -3y = -3 y = 1 so x = 4 and y = 1......or (4,1)

OpenStudy (texaschic101):

we added....you see how the 3y's cancelled each other out...then I solved for x once x was found, sub that back in and solve for y

OpenStudy (anonymous):

thanks alot! now i can try to do the others on here.

OpenStudy (texaschic101):

good to hear :)